### v1

# Cubic NK-SVD: An Algorithm for Designing Parametric Dictionary in Frequency Estimation

Abstract
We propose a novel parametric dictionary learning algorithm for line spectral estimation, applicable in both single measurement vector (SMV) and multiple measurement vectors (MMV) scenarios. This algorithm, termed cubic Newtonized K-SVD (NK-SVD), extends the traditional K-SVD method by incorporating cubic regularization into Newton refinements. The proposed Gauss-Seidel scheme not only enhances the accuracy of frequency estimation over the continuum but also achieves better convergence by incorporating higher-order derivative information. We rigorously establish the convergence of the proposed algorithm within the Block Coordinate Descent (BCD) framework. To the best of our knowledge, this is the first convergence analysis of BCD with a higher-order regularization scheme. Furthermore, this convergence framework can be applied to designing general alternating minimization algorithms with higher-order regularization. Extensive simulations demonstrate that cubic NK-SVD outperforms existing state-of-the-art methods in both SMV and MMV settings.
Index Terms:
Parametric dictionary learning, line spectral estimation, Cubic Newton method, block coordinate descent.
I Introduction
The problem of line spectral estimation (LSE) has been an active research area for decades. The reason is that a variety of applications in signal processing can be recast as LSE, including direction-of-arrival (DOA) estimation [1] and channel estimation in wireless communications [2]. In trying to solve the LSE problem, classical subspace methods [3], e.g., the multiple signal classification (MUSIC) [4] and the estimation of parameters by rotational invariant techniques (ESPRIT) [5], exploit the low-rank structure of the autocorrelation matrix. Under the assumption of Gaussian noise and complete data, these methods achieve good estimation performance with a sufficient number of snapshots. However, their performance deteriorates in challenging scenarios, such as when the number of snapshots is small, the frequencies are closely spaced, the model order is unknown, the observed data are incomplete, or the signal-to-noise ratio (SNR) is low.
Recently, the research on LSE problem has been advanced by the development of methods based on compressed sensing (CS) [6]. In these methods, e.g., $\ell_{1}$-SVD [1], a finite discretized dictionary is used to represent all possible DOA estimates. However, accurate reconstruction of a sparse signal depends on knowing the sparsifying dictionary, which is often impractical to predefine in DOA estimation. Additionally, applying CS to such applications requires discretizing the continuous parameter space into a finite set of grid points. This discretization inevitably introduces errors, as the true parameters may not align with the grid points. This error, known as grid mismatch [7], significantly degrades the performance of reconstruction algorithms. We refer to such methods utilizing finite fixed grids as on-grid methods.
To combat the severity of the grid mismatch, so-called off-grid methods have been proposed [8, 9, 10, 11, 12, 13]. In [8], a super-resolution iterative reweighted algorithm (referred to as SURE-IR) is proposed to address the LSE problem using the majorization-minimization method. However, this method is computationally intensive. In [9], a fast sequential Newtonized Orthogonal Matching Pursuit (NOMP) algorithm is proposed. The study shows that NOMP achieves computational efficiency by incorporating a Newton-based cyclic refinement step into the Orthogonal Matching Pursuit (OMP) framework [14]. Furthermore, [10] introduces a multi-snapshot Newtonized Orthogonal Matching Pursuit (MNOMP) algorithm, which extends NOMP from the single measurement vector (SMV) case to the multiple measurement vectors (MMV) case. In [11, 12, 13], a Bayesian view is taken on the LSE problem. Besides these efforts, another line of research [15, 16, 17] explores the problem of grid mismatch in a more fundamental manner by avoiding discretization and directly utilizing continuous dictionaries, thus leading to the development of gridless techniques. The study in [15] introduces an atomic norm minimization (ANM) approach to address the LSE problem with continuous dictionaries. This ANM approach is further extended to the MMV case in [16]. Additionally, [17] proposes another gridless method based on enhanced matrix completion (EMaC). Despite their theoretical advancements, both the ANM methods [15, 16] and the EMaC method [17] involve semidefinite programming (SDP), which can result in significant computational complexity.
The aforementioned off-grid methods [8, 9, 10, 11, 12, 13] can be considered specialized parametric dictionary learning (DL) methods [18, 19], where each atom is determined by a single frequency parameter. These methods typically employ an alternating minimization strategy between the parametric atoms and their corresponding gains. Additionally, some approaches directly address the LSE problem using parametric DL methods [20, 21]. However, these methods are first-order algorithms and do not incorporate second-order information. Moreover, they lack convergence analysis and utilize a parallel update strategy for all parametric atoms, without leveraging a Gauss-Seidel-like accelerated strategy as used by the K-SVD algorithm [22]. The authors in [9] also emphasize that the decision-feedback mechanism in this alternative strategy is crucial for handling interference among the underlying atoms.
In this paper, we address the LSE problem within a parametric DL framework to mitigate the effect of grid mismatch. Inspired by the classical DL approach [22], we extend the K-SVD algorithm [22] to the frequency estimation problem by incorporating the cubic regularization of Newton’s method [23]. The proposed approach, termed cubic Newtonized K-SVD (cubic NK-SVD) algorithm, is suitable for both SMV and MMV scenarios. Another contribution of our work is the first rigorous establishment of the convergence analysis of Block Coordinate Descent (BCD) [24] using higher-order derivative information. This analytical framework is not only applicable to the analysis of our proposed cubic NK-SVD algorithm but can also be used in the design of general alternating minimization algorithms with higher-order regularization. Extensive numerical results are provided to verify our theoretical findings and demonstrate the advantages of the cubic NK-SVD algorithm in terms of estimation accuracy and computational efficiency.
We use the following notation throughout this paper: $\boldsymbol{A}$ is a matrix; $\boldsymbol{a}$ is a vector; $a$ is a scalar; $\boldsymbol{A}^{T}$ and $\boldsymbol{A}^{H}$ denote transpose and conjugate transpose of $\boldsymbol{A}$, respectively; $\|\boldsymbol{A}\|_{F}$ is the Frobenius norm of $\boldsymbol{A}$; $\|\boldsymbol{a}\|_{p}$ is the $p$-norm of $\boldsymbol{a}$; $\|\boldsymbol{a}\|$ specifically denotes the 2-norm of $\boldsymbol{a}$; $\operatorname{Diag}(\boldsymbol{a})$ denotes a diagonal matrix with the elements of vector $\boldsymbol{a}$ on the diagonal; $\boldsymbol{I}_{N}$ is the $N\times N$ identity matrix; $\boldsymbol{0}_{M\times N}$ is the $M\times N$ all-zeros matrix; The real part of a variable is denoted by ${}\Re{\{\cdot\}}$.
The rest of the paper is organized as follows. In Section II, we formulate the LSE problem as a parametric DL problem. The proposed cubic NK-SVD algorithm is developed in Section III. Section IV offers a convergence analysis based on the BCD framework. Simulation results are provided in Section V. Finally, the paper is concluded in Section VI.
II Problem Formulation
Consider the LSE problem where the observed signal is a summation of a number of complex sinusoids:
$\boldsymbol{y}(t)=\boldsymbol{A}(\boldsymbol{\theta})\boldsymbol{s}(t)+\boldsymbol{e}(t),\quad t=1,\cdots,T,$ (1)
where $\boldsymbol{y}(t)=\left[y_{1}(t),\cdots,y_{N}(t)\right]^{T}$, $\boldsymbol{\theta}=\left[\theta_{1},\cdots,\theta_{K}\right]^{T}$, $\boldsymbol{s}(t)=\left[s_{1}(t),\cdots,s_{K}(t)\right]^{T}$, $\boldsymbol{e}(t)=\left[e_{1}(t),\cdots,e_{N}(t)\right]^{T}$. The matrix $\boldsymbol{A}(\boldsymbol{\theta})=\left[\boldsymbol{a}(\theta_{1}),\cdots,\boldsymbol{a}(\theta_{K})\right]^{T}$ is an array manifold matrix and $\boldsymbol{a}(\theta_{k})$ is called steering vector of the $k$-th source. Defining the unit norm steering vector of frequency $\theta\in[0,2\pi)$, by
$\boldsymbol{a}(\theta)=\dfrac{1}{\sqrt{N}}\left[1,e^{j\theta}\cdots,e^{j(N-1)\theta}\right]^{T},$ (2)
the model (1) can be rewritten in a matrix form as
$\boldsymbol{Y}=\boldsymbol{A}(\boldsymbol{\theta})\boldsymbol{S}+\boldsymbol{E},$ (3)
where $\boldsymbol{Y}=[\boldsymbol{y}(1),\cdots,\boldsymbol{y}(T)]$ is the observation matrix consisting of $T$ observed vectors, $\boldsymbol{S}=[\boldsymbol{s}(1),\cdots,\boldsymbol{s}(T)]$, and $\boldsymbol{E}$ denotes the noise matrix.
The issue also arises in the context of CS reconstruction. When a sensing matrix $\boldsymbol{\Phi}\in\mathbb{C}^{M\times N}$ (with $M\ll N$) is employed, the atoms in our model transform into $\tilde{\boldsymbol{a}}(\theta)=\boldsymbol{\Phi}\boldsymbol{a}(\theta)$. In this setting, the atom $\tilde{\boldsymbol{a}}(\theta)$ resides on the new manifold $\tilde{\boldsymbol{A}}(\boldsymbol{\theta})=\left[\tilde{\boldsymbol{a}}(\theta_{1}),\cdots,\tilde{\boldsymbol{a}}(\theta_{K})\right]$ formed by compressive measurements. The goal of this paper is to estimate the number of sources $K$, the unknown frequencies $\boldsymbol{\theta}$, and gains $\boldsymbol{S}$ given the observed data $\boldsymbol{Y}$ and the sensing matrix $\boldsymbol{\Phi}$.
Now, we start to formulate the LSE problem as a parametric DL problem [22]. The problem as it appears in (3) is a nonlinear parameter estimation problem. To cast this problem as a parametric DL problem, we introduce an overcomplete dictionary $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$ in terms of all possible source locations. Let $\boldsymbol{\tilde{\theta}}=[\tilde{\theta}_{1},\cdots,\tilde{\theta}_{R}]$ be a sampling grid of all source locations of interest. The number of potential source locations $R$ will typically be much greater than the number of sources $K$ or even the number of sensors $N$. We construct a parametric dictionary composed of steering vectors corresponding to each potential source location as its columns: $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})=[\boldsymbol{a}(\tilde{\theta}_{1}),\cdots,\boldsymbol{a}(\tilde{\theta}_{R})]$.
We represent the signal field using the matrix $\boldsymbol{X}=[\boldsymbol{x}(1),\cdots,\boldsymbol{x}(T)]$. In the MMV case, the model is formulated as
$\boldsymbol{Y}=\boldsymbol{D}(\boldsymbol{\tilde{\theta}})\boldsymbol{X}+\boldsymbol{E}.$ (4)
In effect, this overcomplete representation allows us to exchange the LSE problem for the problem of designing the parametric dictionary $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$ that leads to the best possible representations $\boldsymbol{X}$ with strict sparsity constraints.
III Cubic NK-SVD Algorithm
In this section, we introduce the cubic NK-SVD algorithm for designing the parametric dictionary $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$ to estimate the unknown parameter $\boldsymbol{\theta}$ and $\boldsymbol{X}$. It is designed to be an application of the classical DL algorithm K-SVD [22] with Cubic Newton refinements to sidestep the potential basis mismatch from discretizing a continuous parameter space. For purposes of exposition, we first derive our algorithm in the MMV case. The SMV is a special case by simply setting $T=1$. The optimization problem is that of searching the best possible parametric dictionary $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$ for the sparse representation of the training set $\boldsymbol{Y}$
$\min_{\boldsymbol{\tilde{\theta}},\boldsymbol{X}}$ $\ \|\boldsymbol{Y}-\boldsymbol{D}(\boldsymbol{\tilde{\theta}})\boldsymbol{X}\|_{F}^{2}$ (5)
s.t. $\ \ \|{{\boldsymbol{x}}_{i}}\|_{0}\leq{K},\ \forall i=1,\cdots,T,$
where ${{\boldsymbol{x}}_{i}}$ is the $i$-th column of the matrix $\boldsymbol{X}$.
In our algorithm, we minimize the objective in (5) iteratively between sparse coding stage and atom update stage. In the sparse coding stage, we fix $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$ and aim to find the best sparse representation $\boldsymbol{X}$. There are many algorithms for solving the sparse representation problem, e.g., OMP [14] and fast iterative shrinkage-thresholding algorithm (FISTA) [25]. In our algorithm, we choose OMP [14] to solve the sparse representation problem, due to its good trade-off between complexity and representation abilities. We use an error bound $\epsilon$, determined by an estimate of noise power, as the stopping criterion for OMP without prior knowledge of model order $K$ [14]. Specifically, in our implementation, OMP expands the support until the residual error becomes smaller than the bound $\epsilon$. In the atom update stage, we use a Gauss–Seidel-like accelerated scheme in [22], i.e., we update the atoms of $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$ sequentially and allow changing the relevant coefficients. The difference from the K-SVD [22] lies in the specific atomic structure of the atoms $\boldsymbol{a}(\tilde{\theta}_{k})$ in the parametric dictionary [^1] The discussion here is not limited to Fourier atom (2). The extension to general atomic sets is straightforward.. Moreover, updates are only required for the parameters $\boldsymbol{\tilde{\theta}}$ within the dictionary $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$. Assuming that both $\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$ and $\boldsymbol{X}$ are fixed except $\theta_{k}$ and the corresponding coefficients $\boldsymbol{x}_{T}^{k}$, the subproblem of estimating $\theta_{k}$ and $\boldsymbol{x}_{T}^{k}$ can be transformed into minimizing the following objective
$S(\tilde{\theta}_{k},\boldsymbol{x}_{T}^{k})=$ $\|\boldsymbol{Y}-\boldsymbol{D}(\boldsymbol{\tilde{\theta}})\boldsymbol{X}\|_{F}^{2}$ (6)
$=$ $\|\boldsymbol{Y}-\sum_{j=1}^{R}\boldsymbol{a}(\tilde{\theta}_{j})\boldsymbol{x}_{T}^{j}\|_{F}^{2}$
$=$ $\|(\boldsymbol{Y}-\sum_{j\neq k}\boldsymbol{a}(\tilde{\theta}_{j})\boldsymbol{x}_{T}^{j})-\boldsymbol{a}(\tilde{\theta}_{k})\boldsymbol{x}_{T}^{k}\|_{F}^{2}$
$=$ $\|\boldsymbol{E_{k}}-\boldsymbol{a}(\tilde{\theta}_{k})\boldsymbol{x}_{T}^{k}\|_{F}^{2},$
where $\boldsymbol{x}_{T}^{j}$ is the coefficients that correspond to the $j$-th atom $\boldsymbol{a}(\tilde{\theta}_{j})$, i.e., the $j$-th row of the matrix $\boldsymbol{X}$. The matrix $\boldsymbol{E_{k}}$ stands for the error for all the $T$ examples when the $k$-th atom is removed.
Directly optimizing $S(\tilde{\theta}_{k},\boldsymbol{x}_{T}^{k})$ over $\tilde{\theta}_{k}$ and $\boldsymbol{x}_{T}^{k}$ is difficult. Therefore, we adopt a two stage procedure: (1) Coarse estimate stage: In this stage, we obtain a coarse estimate of $\tilde{\theta}_{k}$ by restricting it to a discrete set and using singular value decomposition (SVD) to find the closest rank-1 matrix (in Frobenius norm) that approximates $\boldsymbol{E_{k}}$ [22]; (2) Cubic Newton refinement stage: In this stage, we iteratively refine $\tilde{\theta}_{k}$ and $\boldsymbol{x}_{T}^{k}$ using the cubic regularization of Newton’s method [23]. To enforce the sparsity pattern of $\boldsymbol{x}_{T}^{k}$, we use a similar scheme to that described in [22]. Let $\mathcal{I}_{k}$ denote the indices of the nonzero positions of $\boldsymbol{x}_{T}^{k}$, i.e., $\mathcal{I}_{k}=\{\ell\mid\boldsymbol{x}_{T}^{k}(\ell)\neq 0\}$. Define $\boldsymbol{E_{k}^{R}}=(\boldsymbol{E_{k}})_{\mathcal{I}_{k}}$, where the subscript $\mathcal{I}_{k}$ indicates the restriction of the matrix to the columns with indices in $\mathcal{I}_{k}$. Similarly, $\boldsymbol{x}_{R}^{k}=(\boldsymbol{x}_{T}^{k})_{\mathcal{I}_{k}}$ is obtained by removing the zero entries from the row vector $\boldsymbol{x}_{T}^{k}$. With this notation, the minimization of (6) can be performed directly via SVD, disregarding the parametric structure of the atoms. Applying SVD to the restricted matrix $\boldsymbol{E_{k}^{R}}$, we decompose it as $\boldsymbol{E_{k}^{R}}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{H}$. We define $\boldsymbol{u}_{1}$ as the first column of $\boldsymbol{U}$. Further considering the structure of the atom, the atom corresponding to optimal $\tilde{\theta}_{k}$ should be sufficiently close to $\boldsymbol{u}_{1}$. Using this observation, we can find a coarse estimate of $\tilde{\theta}_{k}$ in the coarse estimate stage.
Coarse estimate stage: We obtain a coarse estimate of $\tilde{\theta}_{k}$ by restricting it to a finite discrete set denoted by ${\boldsymbol{\Omega}}=\{k(2\pi/{\gamma}N):k=0,1,\cdots,({\gamma}N-1)\}$, where $\gamma$ is the oversampling factor relative to the Discrete Fourier Transform (DFT) grid. The output $(\tilde{\theta}_{k})_{c}$ of this stage is the solution to the following optimization problem
$(\tilde{\theta}_{k})_{c}=\arg\min_{\theta\in{\boldsymbol{\Omega}}}|\boldsymbol{a}(\theta)^{H}\boldsymbol{u}_{1}|.$ (7)
Given $\tilde{\theta}_{k}$ is fixed, the optimal coefficients that minimizes $S(\tilde{\theta}_{k},\boldsymbol{x}_{T}^{k})$ is given by
$\hat{\boldsymbol{x}}_{R}^{k}=\boldsymbol{a}(\tilde{\theta}_{k})^{H}\boldsymbol{E_{k}^{R}},$ (8)
where $\hat{\boldsymbol{x}}_{R}^{k}=(\hat{\boldsymbol{x}}_{T}^{k})_{\mathcal{I}_{k}}$. Thus, the corresponding coefficients in set $\mathcal{I}_{k}$ of this stage are $\boldsymbol{a}((\tilde{\theta}_{k})_{c})^{H}\boldsymbol{E_{k}^{R}}$.
Cubic Newton refinement stage: To combat the severity of the grid mismatch in coarse estimate stage, we operate a Newton-based refinement for estimation on the continuum. Let $(\hat{\theta}_{k},\hat{\boldsymbol{x}}_{T}^{k})$ denote the current estimate. Note that the partial Hessian of our objective function $S$ with respect to $\theta$ is Lipschitz continuous, as defined in Assumption 1(ii). Thus, we can try to refine the frequency $\hat{\theta}_{k}$ to $\hat{\theta}_{k}^{\prime}$ using cubic regularization of Newton’s method proposed in [23]. The refinement step for frequency is given by
$\hat{\theta}_{k}^{\prime}\in\arg\min_{\theta}\xi_{\hat{\theta}_{k},\hat{\boldsymbol{x}}^{k}_{T}}^{L(\hat{\boldsymbol{x}}^{k}_{T})}(\theta),$ (9)
where the auxiliary function
$\xi_{\hat{\theta}_{k},\hat{\boldsymbol{x}}^{k}_{T}}^{L(\hat{\boldsymbol{x}}^{k}_{T})}(\theta)=$ $\left\langle{{\nabla_{\theta}}S{\left(\hat{\theta}_{k},\hat{\boldsymbol{x}}^{k}_{T}\right)}},\theta-\theta_{k}\right\rangle$ (10)
$+\frac{1}{2}\left\langle{{\nabla^{2}_{\theta\theta}}S\left(\hat{\theta}_{k},\hat{\boldsymbol{x}}^{k}_{T}\right)}(\theta-\theta_{k}),\theta-\theta_{k}\right\rangle$
$+\frac{{L(\hat{\boldsymbol{x}}^{k}_{T})}}{6}|\theta-\theta_{k}|^{3}$
is an upper second-order approximation for our objective function. $L(\hat{\boldsymbol{x}}^{k}_{T})$ is the Lipschitz continuous moduli of objective function given $\hat{\boldsymbol{x}}^{k}_{T}$. The gain parameter is then updated according to (8). In Section IV, we will provide a detailed analysis of the aforementioned strategy and demonstrate the convergence guarantee of this scheme throught a BCD framework [24].
An event that frequently occurs is is the non-utilization of an atom. Specifically, if the unused atom is $\boldsymbol{a}(\tilde{\theta}_{j})$, it implies that $\mathcal{I}_{j}=\emptyset$. Consequently, updating the atom becomes meaningless since it cannot be modified. Therefore, specific actions must be taken for unused atoms. Our approach involves directly removing these unused atoms, which allows us to retain only the relevant ones. This strategy not only yields a more efficient set of active atoms but also reduces the computational complexity of the algorithm, as the number of useful atoms is considerably smaller than the initial grid resolution $R$. The cubic NK-SVD algorithm [^2] It can be easily extended to compressive scenarios with a compressive manifold $\boldsymbol{\tilde{D}}(\boldsymbol{\tilde{\theta}})=\boldsymbol{\Phi}\boldsymbol{D}(\boldsymbol{\tilde{\theta}})$. is summarized in Algorithm 1.
Algorithm 1 Cubic Newtonized K-SVD (Cubic NK-SVD)
1:  Input: $\boldsymbol{Y}$, $R$, $\gamma$, $\epsilon$.
2:  Initialize: $\widehat{K}=R$, $\boldsymbol{f}=\{k(2\pi/R):k=0,1,\cdots,R-1\}$ and $\boldsymbol{\Omega}=\{k(2\pi/{\gamma}N):k=0,1,\cdots,({\gamma}N-1)\}$.
3:  Output: Updated $\widehat{K}$, $\boldsymbol{f}$, and $\boldsymbol{X}$.
4:  repeat
5:     Sparse Coding Stage:
6:     $\boldsymbol{X}$ is updated using OMP with an error bound $\epsilon$.
7:     Atom Update Stage:
8:     for $k=$ 1 to $R$ do
9:        Compute $\mathcal{I}_{k}=\{\ell\mid{\boldsymbol{x}}_{T}^{k}(\ell)\neq 0\}$.
10:        if $\mathcal{I}_{k}=\emptyset$ then
11:           Eliminate the unused atom $\boldsymbol{a}(\tilde{\theta}_{j})$.
12:           $\widehat{K}=\widehat{K}-1$.
13:        else
14:           Compute error $\boldsymbol{E_{k}}=\boldsymbol{Y}-\sum_{j\neq k}\boldsymbol{a}(\tilde{\theta}_{j}){\boldsymbol{x}}_{T}^{j}$.
15:           Retrict $\boldsymbol{E_{k}}$ by $\boldsymbol{E_{k}^{R}}=(\boldsymbol{E_{k}})_{\mathcal{I}_{k}}$.
16:           Compute first left singular vector $\boldsymbol{u}_{1}$ of $\boldsymbol{E_{k}^{R}}$.
17:           Coarse Estimate: Update $(\tilde{\theta}_{k})_{c}$ within $\boldsymbol{\Omega}$ by (7) and its corresponding gain $({\boldsymbol{x}}_{T}^{k})_{c}$ by (8).
18:           Cubic Newton Refinement: Refine $(\hat{\theta}_{k},\hat{\boldsymbol{x}}_{T}^{k})$ using (9).
19:        end if
20:     end for
21:  until stopping criterion met
Remark 1.
Stopping criterion: Cubic NK-SVD algorithm will be terminated if the difference in function values between two consecutive steps is less than $10^{-3}$, or a maximum of 30 iterations is reached.
IV Convergence
We begin by discussing the estimation of a single frequency and then extend this approach to a mixture of frequencies. Initially, we focus on solving the nonconvex minimization problem involving two blocks of variables, $(x,\boldsymbol{y})\in\mathbb{R}^{1}\times\mathbb{C}^{T}$:
$\min_{{x},\boldsymbol{y}}H({x},\boldsymbol{y}).$ (11)
In line with the approach outlined in [24] and [23], we make the following blanket assumption.
Assumption 1.
(i) Assume that the objective function $H(x,\boldsymbol{y})$ is bounded below, i.e., $\inf_{\mathbb{R}^{1}\times\mathbb{C}^{T}}H\geq H^{*}.$
(ii) For any fixed $\boldsymbol{y}$ the function $x\rightarrow H(x,\boldsymbol{y})$ is $C_{L(\boldsymbol{y})}^{2,2}$, namely the partial Hessian $\nabla^{2}_{xx}H(x,\boldsymbol{y})$ is Lipschitz continuous with moduli $L(\boldsymbol{y})$, that is
$\left\|\nabla^{2}_{xx}H\left(x_{1},\boldsymbol{y}\right)-\nabla^{2}_{xx}H\left(x_{2},\boldsymbol{\boldsymbol{y}}\right)\right\|$ $\leq L(\boldsymbol{y})\left\|x_{1}-x_{2}\right\|,$
$\forall x_{1},x_{2}\in\mathbb{R}^{1}.$
(iii) For any fixed $x$ the function $\boldsymbol{y}\rightarrow H(x,\boldsymbol{y})$ is assumed to be strongly convex (or, more specifically, $\nu$-strongly convex) in $\mathbb{C}^{T}$. In other words, there exists $\nu>0$ such that
${(\nabla^{2}_{\boldsymbol{y}\boldsymbol{y}})_{c}H(x,\boldsymbol{y})}\succeq\frac{\nu}{2}{\boldsymbol{I}_{2T}}\succ{\boldsymbol{0}},$
where we used the notation ${(\nabla^{2}_{\boldsymbol{y}\boldsymbol{y}})_{c}H(\boldsymbol{y})}$ to refer to the complex form of the partial Hessian matrix of $H(x,\boldsymbol{y})$. For the reader’s convenience, the related definitions are summarized in the Appendix B (see Appendices in [26] for more details).
(iv) There exists $\lambda^{-},\lambda^{+}$ such that
$\inf\left\{L\left({\boldsymbol{y}}^{k}\right):k\in\mathbb{N}\right\}\geq\lambda^{-},$ (12)
$\sup\left\{L\left({\boldsymbol{y}}^{k}\right):k\in\mathbb{N}\right\}\leq\lambda^{+}.$ (13)
A few words on Assumption 1 are now in order.
Remark 2.
(i) The partial Lipschitz properties required in Assumption 1(ii) are at the heart of the cubic NK-SVD algorithm which is designed to fully exploit the Hessian-Lipschitz property of the problem at hand.
(ii) In our problem, the goal is to estimate unknown frequencies and corresponding gains. To be specific, the objective function $H(x,\boldsymbol{y})$ is
$H(x,{\boldsymbol{y}})=\|\boldsymbol{E}-\boldsymbol{a}(x){\boldsymbol{y}^{H}}\|_{F}^{2},$ (14)
where $\boldsymbol{a}(x)$ is the steering vector of frequency $x\in\mathbb{R}^{1}$. It is easy to verify that $H(x,{\boldsymbol{y}})$ in (14) satisfies the above Assumption 1 (See Appendix C).
As outlined in Algorithm 1, the cubic NK-SVD algorithm can be viewed as a Gauss-Seidel iteration scheme, popularized in modern era under the name alternating minimization. That is, starting with some given initial point $(x^{0},{\boldsymbol{y}}^{0})$, we generate a sequence $\{(x^{k},{\boldsymbol{y}}^{k})\}_{k\in\mathbb{N}}$ via the scheme
$x^{k+1}\in\arg\min_{x}\xi_{x^{k},{\boldsymbol{y}}^{k}}^{{L({\boldsymbol{y}}^{k})}}(x),$ (15)
${\boldsymbol{y}}^{k+1}\in\arg\min_{\boldsymbol{y}}H\left(x^{k+1},\boldsymbol{y}\right),$ (16)
where the auxiliary function
$\xi_{x^{k},{\boldsymbol{y}}^{k}}^{{L(\boldsymbol{y}^{k})}}(x)=$ $\left\langle\nabla_{x}H\left(x^{k},{\boldsymbol{y}}^{k}\right),x-x^{k}\right\rangle$ (17)
$+\frac{1}{2}\left\langle\nabla^{2}_{xx}H\left(x^{k},{\boldsymbol{y}}^{k}\right)(x-x^{k}),x-x^{k}\right\rangle$
$+\frac{{L(\boldsymbol{y}^{k})}}{6}|x-x^{k}|^{3}$
will be an upper second-order approximation for our objective function under the Assumption 1(ii):
$H(x,{\boldsymbol{y}}^{k})\leq\xi_{x^{k},{\boldsymbol{y}}^{k}}^{{L(\boldsymbol{y}^{k})}}(x),\quad\forall x\in\mathbb{R}^{1}.$
This is exactly the approach we analyze in this paper. This scheme is called cubic regularization of Newton’s method [23]. Note that subproblem (15) is non-convex and may have local minima. However, our approach remains practical since the problem involves minimization over a single variable. Specifically, subproblem (15) can be reformulated as solving the following nonlinear equation:
$\nabla_{x}H\left(x^{k},{\boldsymbol{y}}^{k}\right)+\nabla^{2}_{xx}H\left(x^{k},{\boldsymbol{y}}^{k}\right)(x-x^{k})$ (18)
$+\frac{1}{2}{L(\boldsymbol{y}^{k})}|x-x^{k}|\cdot(x-x^{k})=0.$
Due to the strong convexity assumption of $h$ with respect to $\boldsymbol{y}$, the minimum of (16) in each step is uniquely attained
$\boldsymbol{y}^{k+1}=\boldsymbol{E}^{H}\boldsymbol{a}(x^{k+1}).$ (19)
Convergence results for the Gauss-Seidel method, also known as the BCD approach, can be found in several studies [27, 24]. In the light of general methodology in [24], we characterize the convergence results for the cubic NK-SVD algorithm.
We first establish some useful properties for the cubic NK-SVD algorithm under our Assumption 1. For convenience, we will use the notation
$\boldsymbol{z}^{k}:=(x^{k},\boldsymbol{y}^{k}),\quad\forall k\geq 0.$
Lemma 1.
Suppose that Assumption 1 hold. Let $\{\boldsymbol{z}^{k}\}_{k\in\mathbb{N}}$ be a sequence iteratively generated by (15) and (16). The following assertions hold.
(i) The sequence $\{H(\boldsymbol{z}^{k})\}_{k\in\mathbb{N}}$ is nonincreasing, and in particular, for any $k\geq 0$, we have
$\frac{\lambda^{-}}{12}\left|x^{k+1}-x^{k}\right|^{3}+\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|^{2}\leq H\left(\boldsymbol{z}^{k}\right)-H\left(\boldsymbol{z}^{k+1}\right).$ (20)
(ii) We have
$\sum_{k=0}^{\infty}\left|x^{k+1}-x^{k}\right|^{{3}}+\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|_{3}^{{3}}$ (21)
$=\sum_{k=0}^{\infty}\left\|\boldsymbol{z}^{k+1}-\boldsymbol{z}^{k}\right\|_{3}^{{3}}<\infty,$
and hence $\lim_{k\rightarrow\infty}\left\|\boldsymbol{z}^{k+1}-\boldsymbol{z}^{k}\right\|=0$.
Proof.
See Appendix A-B. ∎
Notice that the function $\boldsymbol{y}\rightarrow H(x,\boldsymbol{y})$ (with $x$ is fixed) is a real-valued function of the complex variable $\boldsymbol{y}\in\mathbb{C}^{T}$. It is straightforward to verify that any other nonconstant real-valued function $g(\boldsymbol{z})$ is not holomorphic in $\boldsymbol{z}$; therefore, it need not be differentiable in the traditional sense of complex differentiation. However, in many instances, we are only interested in determining the locations of the stationary points of $g(\boldsymbol{z})$. For these cases, it is sufficient to rely on a different definition of differentiation based on what is sometimes known as the Wirtinger calculus [26] (see Appendix B).
To illustrate the benefit of our approach, we now show how iterative scheme (15) and (16), which is the heart of the cubic NK-SVD algorithm, can be applied to efficiently solve the LSE problem with objective function $H(x,\boldsymbol{y})$ defined in (14).
Lemma 2.
Suppose that Assumptions 1 hold. Let $\{\boldsymbol{z}^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (15) and (16) which is assumed to be bounded, i.e., there exists $\sigma$ such that $\|\boldsymbol{z}^{k}\|\leq\sigma,\forall k\in\mathbb{N}$. For each positive integer $k$, we have
$\|\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k+1})\|\leq\lambda^{+}\|x^{k+1}-x^{k}\|^{2}+2{\sigma}\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\|,$ (22)
and Wirtinger derivative
$\|\nabla_{\boldsymbol{y}}H(x^{k+1},\boldsymbol{y}^{k+1})\|={\boldsymbol{0}}.$ (23)
Then,
$\|\nabla_{\boldsymbol{z}}H(\boldsymbol{z}^{k+1})\|$ (24)
$\leq\|\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k+1})\|+\|\nabla_{\boldsymbol{y}}H(x^{k+1},\boldsymbol{y}^{k+1})\|$
$\leq{\rho_{2}}\|\boldsymbol{z}^{k+1}-\boldsymbol{z}^{k}\|,$
where $\rho_{2}=\max\{\lambda^{+},2\sigma\}$.
Proof.
See Appendix A-C. ∎
In the following, we will summarize several properties of the limit point set. Let $\{\boldsymbol{z}^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by the cubic NK-SVD algorithm from a starting point $\boldsymbol{z}^{0}$. The set of all limit points is denoted by $\omega(\boldsymbol{z}^{0})$, i.e.,
$\omega(\boldsymbol{z}^{0})=\{\bar{\boldsymbol{z}}\in\mathbb{R}^{1}\times\mathbb{C}^{T}:\exists\ \text{an increasing sequence of integers}$
$\{k_{l}\}_{l\in\mathbb{N}},\ \text{sucn that}\ \boldsymbol{z}^{k_{l}}\rightarrow\bar{\boldsymbol{z}}\ as\ l\rightarrow\infty\}.$
Theorem 1.
Suppose that Assumption 1 hold. Let $\{\boldsymbol{z}^{k}\}_{k\in\mathbb{N}}$ be a sequence generated from the cubic NK-SVD algorithm which is assumed to be bounded. The following assertions hold.
(i) $\emptyset\neq\omega(\boldsymbol{z}^{0})\in crit\,H:=\{\boldsymbol{z}:\nabla_{\boldsymbol{z}}H(\boldsymbol{z})=0\}$.
(ii) We have
$\lim_{k\rightarrow\infty}{dist}(\boldsymbol{z}^{k},\omega(\boldsymbol{z}^{0}))=0,$ (25)
where ${dist}(\boldsymbol{z}^{k},\omega(\boldsymbol{z}^{0}))$ denotes the distance from $\boldsymbol{z}^{k}$ to $\omega(\boldsymbol{z}^{0})$.
Proof.
(i) Let $\boldsymbol{z}^{*}=(x^{*},\boldsymbol{y}^{*})$ be the limit point of $\{\boldsymbol{z}^{k}\}_{k\in\mathbb{N}}=(x^{k},\boldsymbol{y}^{k})_{k\in\mathbb{N}}$. This means that there is a subsequence $(x^{k_{q}},\boldsymbol{y}^{k_{q}})_{q\in\mathbb{N}}$ such that $(x^{k_{q}},\boldsymbol{y}^{k_{q}})\rightarrow(x^{*},\boldsymbol{y}^{*})$ as $q\rightarrow\infty$. Since $H$ is continuous (See Assumption 1(ii), (iii)), we obtian that
$\lim_{q\rightarrow\infty}H(x^{k_{q}},\boldsymbol{y}^{k_{q}})=H(x^{*},\boldsymbol{y}^{*}).$
On the other hand we know from Lemma 1(ii) and 2 that $\nabla H(\boldsymbol{z})\rightarrow\boldsymbol{0}$ as $k\rightarrow\infty$. This proves that $(x^{*},\boldsymbol{y}^{*})$ is a critical point of $H$.
(ii) This item follows as an elementary consequence of the definition of limit points. ∎
Remark 3.
(i) The simple structure of the cubic NK-SVD algorithm allows for its extension to a more general setting involving $R$ frequencies (i.e., $2R$ blocks), for which Theorem 1 remains valid. The derivations are not detailed here since they are similar to the development for the single frequency case.
(ii) We observe that the sparse coding stage can accelerate algorithm convergence. By selecting useful atoms, the number of variables requiring iteration is reduced. This reduction is beneficial because, in practical scenarios, the number of frequencies to be recovered is significantly smaller than the initial grid size $R$. However, further theoretical analysis is needed and remains a focus for our future research.
(iii) Our convergence results are not limited to the analysis of the proposed cubic NK-SVD algorithm but can also be applied to general alternating minimization algorithms with higher-order regularization [28, 29]. However, this is beyond the scope of this paper, and we consider further analysis as future work.
Figure 1: Norm of residuals versus the number of iterations in 10 Monte Carlo trials.
Figure 2: RSNRs of respective algorithms. (a) RSNRs vs. $M$, $K=7$ and PSNR=10 dB. (b) RSNRs vs. $K$, $M=32$ and PSNR=10 dB.
Figure 3: Success rates of respective algorithms vs. $K$, $M=32$ and PSNR=10 dB.
Figure 4: Frequency estimation using different algorithms when $M=24$, $K=3$ and PSNR=0 dB. (a) Ground truth. (b) Cubic NK-SVD. (c) NOMP. (d) ANM. (e) EMaC. (f) OMP.
V Simulation Results
We now carry out experiments to illustrate the performance of our proposed cubic NK-SVD algorithm for the DOA estimation. We consider a mixture of $K$ sinusoids of length $N=64$. The true frequencies $\{\theta_{k},k=1,\cdots,K\}$ are uniformly generated over $[0,2\pi)$. The true coefficients in $\boldsymbol{s}(t)$ are generated i.i.d. with uniform random phase on $[0,2\pi)$ and amplitudes drawn from a normal density of mean 10 and variance 3. We use this specification to ensure that all components can be distinguished from noise. The observation quality is measured by the peak-signal-to-noise ratio (PSNR) which is defined as $\text{PSNR}=10\log_{10}(1/\sigma^{2})$, where $\sigma^{2}$ denotes the noise variance. We perform 300 Monte Carlo trials for each scenarios. We introduce three metrics to evaluate the recovery performance of respective algorithms, namely, the reconstruction signal-to-noise ratio (RSNR) [8], $\beta(\boldsymbol{\theta},\boldsymbol{\hat{\theta}})$ [12] and success rate. The RSNR is defined as:
$\operatorname{RSNR}=20\log_{10}\left(\frac{\left\|\boldsymbol{A}(\boldsymbol{\theta})\boldsymbol{s}(t)\right\|_{2}}{\left\|\boldsymbol{A}(\boldsymbol{\theta})\boldsymbol{s}(t)-\boldsymbol{A}(\boldsymbol{\hat{\theta}})\boldsymbol{\hat{s}}(t)\right\|_{2}}\right),$ (26)
which measures the reconstruction accuracy of the original signal. The second metric evaluates the reconstruction performance of frequencies and is defined as:
$\beta(\boldsymbol{\theta},\boldsymbol{\hat{\theta}})=\frac{1}{K}\sum_{k=1}^{K}(\min_{\hat{\theta}\in\boldsymbol{{\hat{\theta}}}}d(\hat{\theta},\theta_{k}))^{2},$ (27)
where $d(\hat{\theta},\theta_{k})$ represents the distance between the estimated and true frequencies. The success rate is the proportion of Monte Carlo trials in which the frequency vector $\boldsymbol{\theta}$ is successfully recovered. A trial is considered successful if the model order $K$ is estimated correctly and $\beta(\boldsymbol{\theta},\boldsymbol{\hat{\theta}})<{10^{-3}}$.
We compare our proposed algorithm with other existing state-of-the-art (SOTA) super-resolution methods, namely, subspace method: MUSIC [4]; On-Grid method: OMP [14], $\ell_{1}$-SVD [1]; Off-Grid methods: NOMP [^3] Available at https://bitbucket.org/wcslspectralestimation/continuous-frequency-estimation/src/NOMP/ [9], MNOMP [^4] Available at https://github.com/RiverZhu/MNOMP-code [10]; Gridless methods: ANM [^5] Available at https://github.com/badrinarayan/astlinespec [15], ANM-MMV [^6] Available at https://users.ece.cmu.edu/~yuejiec/papers/atomic_mmv.zip [16], EMaC [^7] Available at https://users.ece.cmu.edu/~yuejiec/papers/EMaC_codes.zip [17]. The Cramér-Rao bound (CRB) [30] is presented as a benchmark that provides a lower bound for any unbiased estimator.
For the cubic NK-SVD algorithm, we set the intial grid resolution $R=N$, the oversampling factor $\gamma=10$, and the error bound $\epsilon=\sqrt{M}\sigma$. MUSIC [4] operates by constructing an estimate of the autocorrelation matrix of the observed data vector. In MMV sccenarios with complete measurements, we use the total $T$ snapshots to generate the autocorrelation matrix. MUSIC and $\ell_{1}$-SVD methods require to prespecify the initial grid points. In our experiments, the initial number of grid points are set to be 360 to guarantee the recovery accuracy. We feed into MUSIC and $\ell_{1}$-SVD the true model order $K$ to pick out the top $K$ peaks, which is implemented using the MATLAB routine findpeaks. The OMP solution is obtained using a grid of size $1282N=$ 128. We have observed that a finer grid does not improve performance. The true model order $K$ is used as a stopping criterion for the OMP algorithm. In NOMP and MNOMP, the oversampling factor $\gamma$ is set to 4. The number of Newton steps $R_{s}$ is set to 1, while the number of cyclic refinements $R_{c}$ is set to 3 as suggested in [9, 10]. The ANM, ANM-MMV and EMaC are solved using SDP approach. The solutions to the primal problems of these gridless methods directly provide an estimate of the signal vector $\boldsymbol{A}(\boldsymbol{\theta})\boldsymbol{s}$. The frequencies are recovered using the root-MUSIC algorithm [31], assuming the true model order $K$ is given. The coefficients are then estimated through least-squares.
V-A Frequency Estimation for SMV Model
We examine the behavior of various algorithms under SMV scenarios with $T=1$. Note that the performance of many existing algorithms depends on minimum frequency separation $\min_{k\neq l}|\theta_{k}-\theta_{l}|$ between two sinusoids in the mixture, denoted as $\Delta\theta_{\text{min}}$. According to [32], if $\Delta\theta_{\text{min}}$ exceeds $4\pi/N$, the ANM approach is known to exactly recover the frequencies in a noise-free scenario. We first consider an experimental setup where $\Delta\theta_{\text{min}}$ is ensured to be greater than $4\pi/N$. The sensing matrix $\boldsymbol{\Phi}$ is configured as a random subsampling matrix, meaning the measurements are obtained by randomly selecting $M$ entries from $N$ elements of the signal.
We first evaluate the convergence property of the cubic NK-SVD algorithm. We set $M=32$, $K=3$, and PSNR=20 dB. We conduct 10 Monte Carlo trials for this scenario. Fig. 1 presents the norm of residuals in (5) versus the number of iterations. The results show that the residuals of all trials tend to converge, corroborating the result in Theorem 1.
In Fig. 2a, we plot the average RSNRs of respective algorithms as a function of the number of measurement, $M$, with $K=7$ and PSNR=10 dB. The proposed method demonstrates superior performance compared to the other four methods, particularly outperforming the NOMP method when $M$ is small. As $M$ increases, the recovery performance of the proposed method approaches the CRB. In Fig. 2b, we present the average RSNRs of respective algorithms vs. the number of frequencies, $K$, with $M=32$ and PSNR=10 dB. It can be observed that our proposed method outperforms other methods for a large number of frequencies. For instance, when $K=10$, our algorithm achieves an approximately 3 dB gain in RSNR compared to the NOMP and ANM approaches. Notably, when $K$ is small, the PSNR of the proposed algorithm exceeds the CRB because the cubic NK-SVD algorithm is biased.
In Fig. 3, we present the success rates of respective algorithms vs. the number of frequencies, $K$, where we set $M=32$ and PSNR=10 dB. The proposed method and the NOMP approach demonstrate significantly higher success rates compared to the gridless method. This superiority is due to the gridless methods’ sensitivity to noise, resulting in unstable performance at lower PSNR levels. Additionally, the OMP method performs the worst, further emphasizing the importance of overcoming grid mismatch.
Let $M=24$, $K=3$, and PSNR=0 dB. In Fig. 4a, we present a spectrally sparse ground truth scene. Figs. 4b4f show the estimated frequencies on a unit circle for different methods. It can be observed that the ANM and EMaC approaches exhibit poor frequency recovery in low PSNR scenarios. In contrast, the proposed method and NOMP approach demonstrate more accurate recovery. Notably, the performance of OMP is hindered by the grid mismatch, preventing it from achieving super-resolution.
Fig. 5 presents the average running times of the respective algorithms, where we set $M=24$ and PSNR=10 dB. The results indicate that NOMP is the most computationally efficient algorithm. The computational efficiency of cubic NK-SVD is notably higher than that of the gridless method [15, 17] and is comparable to NOMP. The performance of cubic NK-SVD is affected by the initial grid resolution $R$ and the oversampling factor $\gamma$. In practical applications, reducing the grid resolution can improve computational efficiency without compromising accuracy. Moreover, in MMV scenarios, incorporating dimension reduction techniques [33] during the sparse coding stage can exploit the joint sparsity of the data while further reducing computational complexity. This approach will be explored in our future work.
Figure 5: Average running times (sec) of respective algorithms when $M=24$ and PSNR=10 dB.
Figure 6: RSNRs of respective algorithms vs. $\mu$, $M=32$, $K=2$ and PSNR=10 dB.
Finally, we evaluate the ability of various algorithms to recover closely-spaced frequencies. We set $\Delta\theta_{\text{min}}=\mu\times 2\pi/N$, where $\mu$ is the frequency spacing coefficient. Fig. 6 illustrates the average RSNRs of different algorithms under different $\mu$ values, with $M=32$, $K=2$, and PSNR=10 dB. It can be observed that cubic NK-SVD significantly outperforms other methods in this challenging scenario. Specifically, when $\mu<1$, the performance of NOMP becomes unstable and degrades significantly, while cubic NK-SVD maintains a notable advantage. The underlying reason is that NOMP, being a greedy method, may approximate two closely-spaced atoms with a single representative atom, especially when $\mu$ is small. In contrast, cubic NK-SVD utilizes a predefined dictionary, allowing it to effectively leverage the interactions between atoms during iterations and achieve super-resolution. Gridless methods struggle to accurately distinguish frequency components in closely-spaced scenarios. The OMP method exhibits the poorest performance in these scenarios, primarily due to grid mismatch. These experimental results highlight the superiority of cubic NK-SVD in handling closely-spaced frequency recovery tasks.
V-B Frequency Estimation for MMV Model
We investigate the performance of cubic NK-SVD in MMV scenarios. In Fig. 7, we plot the angular spectra of respective algorithms. We compare the results obtained via cubic NK-SVD against those obtained via MNOMP [10], ANM-MMV [16], $\ell_{1}$-SVD [1], and MUSIC [4]. As shown in Fig. 7, all methods successfully estimate the frequencies when $M=64$, $\boldsymbol{\Phi}={\boldsymbol{I}}$, $T=48$, $K=5$, and PSNR=20 dB. To better illustrate performance, we plot the $\beta(\boldsymbol{\theta},\boldsymbol{\hat{\theta}})$ of respective algorithms vs. PSNR. Fig. 8 demonstrates that our approach significantly outperforms grid-based methods such as $\ell_{1}$-SVD and MUSIC in frequency recovery. It also performs comparably to ANM-MMV and MNOMP methods. This further illustrates the importance of adjusting the grid in the continuous domain to overcome the grid mismatch problem.
Figure 7: Angular spectra obtained using different algorithms when $M=64$, $K=5$ and PSNR=20 dB.
Figure 8: $\beta(\boldsymbol{\theta},\boldsymbol{\hat{\theta}})$ of respective algorithms vs. PSNR, $M=64$, $T=48$, and $K=5$.
VI Conclusion
In this paper, we addressed the LSE problem by leveraging the framework of parametric DL. We integrated the Cubic Newton refinement into the classic K-SVD algorithm, resulting in a novel parametric DL algorithm. This combination is natural and is thoroughly explained in our convergence analysis. By building upon the BCD framework, we establish a rigorous convergence guarantee for our proposed cubic NK-SVD algorithm. Furthermore, this analytical framework can be extended to design general alternating minimization algorithms with higher-order regularization. Future work may explore the extension of our techniques to other BCD algorithms that utilize higher-order derivative information. The proposed algorithm can be applied to both SMV and MMV scenarios, and an extension for its application in compressive scenarios is also provided. Simulation results demonstrate that the proposed algorithm outperforms SOTA algorithms in various scenarios for DOA estimation.
Appendix A Convergence Analysis
A-A Basic Properties
We first establish some basic properties of the cubic NK-SVD algorithm under our Assumption 1, see e.g., [23, 26].
Lemma 3 ([23]).
Let $h:\mathbb{R}^{1}\rightarrow\mathbb{R}$ be a twice differentiable function with Hessian assumed $L_{h}$-Lipschitz continuous. Then,
$\left|h^{\prime}(x_{2})-h^{\prime}(x_{1})-h^{\prime\prime}(x_{1})(x_{2}-x_{1})\right|\leq\frac{1}{2}L_{h}|x_{2}-x_{1}|^{2},$ (28)
$\left|h(x_{2})-h(x_{1})-\left\langle h^{\prime}(x_{1}),x_{2}-x_{1}\right\rangle\right.$ (29)
$\left.-\frac{1}{2}\left\langle h^{\prime\prime}(x_{1})(x_{2}-x_{1}),x_{2}-x_{1}\right\rangle\right|$ $\leq\frac{L_{h}}{6}|x_{2}-x_{1}|^{3}.$
Define a modified Newton step using the following cubic regularization of quadratic approximation of function
$T_{L_{h}}(x)\in\arg\min_{\tilde{x}}$ $\left[\left\langle h^{\prime}(x),{\tilde{x}}-x\right\rangle+\frac{1}{2}\left\langle h^{\prime\prime}(x)({\tilde{x}}-x),{\tilde{x}}-x\right\rangle\right.$ (30)
$\left.+\frac{L_{h}}{6}|{\tilde{x}}-x|^{3}\right].$
Note that point $T_{L_{h}}(x)$ satisfies the following system of nonlinear equations:
$h^{\prime}(x)+h^{\prime\prime}(x)({\tilde{x}}-x)+\frac{1}{2}L_{h}|{\tilde{x}}-x|\cdot({\tilde{x}}-x)=0.$ (31)
Define
$\bar{h}_{L_{h}}(x)=\min_{\tilde{x}}$ $\left[h(x)+\left\langle h^{\prime}(x),{\tilde{x}}-x\right\rangle\right.$ (32)
$\left.+\frac{1}{2}\left\langle h^{\prime\prime}(x)({\tilde{x}}-x),{\tilde{x}}-x\right\rangle+\frac{L_{h}}{6}|{\tilde{x}}-x|^{3}\right].$
Lemma 4 ([23]).
For any $x\in\mathbb{R}^{1}$ we have
$\bar{h}_{L_{h}}(x)\leq\min_{\tilde{x}}\left[h(\tilde{x})+\frac{L_{h}}{3}|\tilde{x}-x|^{3}\right],$ (33)
$h(x)-\bar{h}_{L_{h}}(x)\geq\frac{L_{h}}{12}|T_{L_{h}}(x)-x|^{3}.$ (34)
Moreover, we have
$h(T_{L_{h}}(x))\leq\bar{h}_{L_{h}}(x).$ (35)
Lemma 5 ([26]).
Let $f:\mathbb{C}^{T}\rightarrow\mathbb{R}$ be a continuously differentiable function that is $\nu$-strongly convex on $\mathbb{C}^{T}$ with $\nabla f(\boldsymbol{y}^{*})=0$. Then,
$f(\boldsymbol{y})\geq f(\boldsymbol{y}^{*})+\frac{\nu}{2}\|\boldsymbol{y}-\boldsymbol{y}^{*}\|^{2},\quad\forall\boldsymbol{y}\in\mathbb{C}^{T}.$ (36)
Lemma 6 ([23]).
$T_{L_{h}}(x)$ satisfies
$|h^{\prime}(T_{L_{h}}(x))|\leq L_{h}|T_{L_{h}}(x)-x|^{2}.$ (37)
A-B Proof of Lemma 1
Proof.
(i) Fix $k\geq 0$. Under our Assumption 1, the functions $x\rightarrow H(x,\boldsymbol{y})$ (with $\boldsymbol{y}$ is fixed) and $\boldsymbol{y}\rightarrow H(x,\boldsymbol{y})$ (with $x$ is fixed) are twice differentiable function with Hessian assumed $L_{h}$-Lipschitz continuous and $\nu$-strongly convex function, respectively. Using the iterative steps (15) and (16), applying Lemma 4 and 5, first with $h(\cdot):=H\left(\cdot,\boldsymbol{y}^{k}\right)$ and $L_{h}=L(\boldsymbol{y}^{k})$, and secondly with $f(\cdot):=H\left(x^{k+1},\cdot\right)$, we obtain successively
$H(x^{k},\boldsymbol{y}^{k})-H(x^{k+1},\boldsymbol{y}^{k})\geq\frac{L(\boldsymbol{y}^{k})}{12}|x^{k+1}-x^{k}|^{3},$ (38)
and
$H(x^{k+1},\boldsymbol{y}^{k})-H(x^{k+1},\boldsymbol{y}^{k+1})\geq\frac{{\nu}}{2}\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\|^{2}.$ (39)
Adding the above two inequalities, we thus obtain for all $k\geq 0$,
$H(\boldsymbol{z}^{k})-H(\boldsymbol{z}^{k+1})\geq$ $\frac{L(\boldsymbol{y}^{k})}{12}|x^{k+1}-x^{k}|^{3}$ (40)
$+\frac{{\nu}}{2}\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\|^{2}.$
From (40) it follows that the sequence $\{H(\boldsymbol{z}^{k})\}_{k\in\mathbb{N}}$ is nonincreasing, and since $H$ is assumed to be bounded below ((see Assumption 1(i))), it converges to some real number $\underline{H}$. Moreover, using the facts that $L(\boldsymbol{y}^{k})\geq\lambda^{-}>0$ ((see Assumption 1(iv))), we get for all $k\geq 0$:
$H(\boldsymbol{z}^{k})-H(\boldsymbol{z}^{k+1})\geq$ $\frac{\lambda^{-}}{12}|x^{k+1}-x^{k}|^{3}$ (41)
$+\frac{{\nu}}{2}\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\|^{2},$
and assertion (i) is proved.
(ii) Let $N$ be a positive integer. Summing (41) from $k=0$ to $N-1$ we also get
$\sum_{k=0}^{N-1}\frac{\lambda^{-}}{12}\left|x^{k+1}-x^{k}\right|^{3}+\frac{{\nu}}{2}\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|^{2}$ (42)
$\leq\left(H\left(\boldsymbol{z}^{0}\right)-H\left(\boldsymbol{z}^{N}\right)\right)$
$\leq\left(H\left(\boldsymbol{z}^{0}\right)-\underline{H}\right).$
Taking the limit as $N\rightarrow\infty$ we abtain
$\sum_{k=0}^{\infty}\frac{\lambda^{-}}{12}\left|x^{k+1}-x^{k}\right|^{3}+\frac{{\nu}}{2}\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|^{2}$ (43)
$\leq\left(H\left(\boldsymbol{z}^{0}\right)-H\left(\boldsymbol{z}^{N}\right)\right)$
$\leq\left(H\left(\boldsymbol{z}^{0}\right)-\underline{H}\right).$
and hence
$\lim_{k\rightarrow\infty}\left|x^{k+1}-x^{k}\right|=0,$ (44)
$\lim_{k\rightarrow\infty}\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|=0.$ (45)
Therefore, when $k$ is sufficiently large, both $\left|x^{k+1}-x^{k}\right|$ and $\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|$ are less than or equal to 1. Without loss of generality, we assume $\left|x^{k+1}-x^{k}\right|\leq 1$, $\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|\leq 1$ hold for all $k\geq 0$. We thus obtain
$\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|_{3}^{3}\leq\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|^{2},\quad\forall k\geq 0.$ (46)
Combining (43) and (46) yields the following
$\sum_{k=0}^{\infty}\rho_{1}\left|x^{k+1}-x^{k}\right|^{3}+\rho_{1}\left\|\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k}\right\|_{3}^{3}$ (47)
$=\sum_{k=0}^{\infty}\rho_{1}\left\|\boldsymbol{z}^{k+1}-\boldsymbol{z}^{k}\right\|_{3}^{3}\leq\left(H\left(\boldsymbol{z}^{0}\right)-\underline{H}\right),$
where $\rho_{1}=\min\{\frac{\lambda^{-}}{12},\frac{\nu}{2}\}$. We obtain the desired assertion (ii). ∎
A-C Proof of Lemma 2
Proof.
Let $k$ be a positive integer. In view of inequality (37), we have
$\left\|\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k})\right\|\leq L(\boldsymbol{y}^{k})\left|x^{k+1}-x^{k}\right|^{2}.$ (48)
Writing down the optimality condition for (16) yields
$\|\nabla_{\boldsymbol{y}}H(x^{k+1},\boldsymbol{y}^{k+1})\|={\boldsymbol{0}}.$ (49)
Note that $H(x,\boldsymbol{y})$ is defined in (14), we have
$\|\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k+1})-\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k})\|$ (50)
$=\left|2\Re\left\{(\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k})^{H}E^{H}\frac{\partial\boldsymbol{a}(x^{k+1})}{\partial x}\right\}\right|$
$=\left|2\Re\left\{(\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k})^{H}\boldsymbol{y}^{k+1}\right\}\right|.$
Then,
$\left\|\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k+1})\right\|\leq$ $\left\|\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k})\right\|$ (51)
$+2\left\|(\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k})\right\|\left\|\boldsymbol{y}^{k+1}\right\|.$
Since we assumed that $\{\boldsymbol{z}^{k}\}_{k\in\mathbb{N}}$ is bounded, we get that
$\left\|\boldsymbol{y}^{k}\right\|\leq\left\|\boldsymbol{z}^{k}\right\|\leq\sigma,\quad\forall k\in\mathbb{N}.$ (52)
The moduli $L(\boldsymbol{y}^{k})$ being bounded from above by $\lambda^{+}$ (see Assumption 1(iv)). In view of inequality (21), we get that when $k$ is sufficiently large, $\left|x^{k+1}-x^{k}\right|$ is less than or equal to 1. Without loss of generality, we assume $\left|x^{k+1}-x^{k}\right|\leq 1,\forall k\geq 0$. Summing up these inequalities, we get
$\left\|\nabla_{x}H(x^{k+1},\boldsymbol{y}^{k+1})\right\|\leq$ $\lambda^{+}\left\|x^{k+1}-x^{k}\right\|^{2}$ (53)
$+2\sigma\left\|(\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k})\right\|$
$\leq$ $\lambda^{+}\left\|x^{k+1}-x^{k}\right\|$
$+2\sigma\left\|(\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k})\right\|$
$\leq$ $\rho_{2}\left\|x^{k+1}-x^{k}\right\|$
$+\rho_{2}\left\|(\boldsymbol{y}^{k+1}-\boldsymbol{y}^{k})\right\|$
$\leq$ $\rho_{2}\left\|\boldsymbol{z}^{k+1}-\boldsymbol{z}^{k}\right\|.$
which is the desired result in (22). Then, in view of inequality (49) and (53), we get the estimate (50). This completes the proof. ∎
Appendix B Wirtinger Calculus in Complex Domain [26]
B-A Complex Gradient Vectors
The Wirtinger partial derivative of $g(\boldsymbol{z})$ with respect to $\boldsymbol{z}\in\mathbb{C}^{T}$ and $\boldsymbol{z}^{*}\in\mathbb{C}^{T}$ are defined as
$\nabla_{\boldsymbol{z}}g(\boldsymbol{{z}}):=\left[\frac{\partial g}{\partial z_{1}},\frac{\partial g}{\partial z_{2}},\cdots\frac{\partial g}{\partial z_{T}}\right]^{T},$ (54)
$\nabla_{\boldsymbol{z}^{*}}g(\boldsymbol{z}):=\left[\frac{\partial g}{\partial z_{1}^{*}},\frac{\partial g}{\partial z_{2}^{*}},\cdots\frac{\partial g}{\partial z_{T}^{*}}\right]^{T},$ (55)
where we express $g({z})$ as a function of its real part $x$ and imaginary part $y$ and define
$\frac{\partial g(z)}{\partial z}:=\frac{1}{2}\left\{\frac{\partial g(x,y)}{\partial x}-j\frac{\partial g(x,y)}{\partial y}\right\},$ (56)
$\frac{\partial g(z)}{\partial z^{*}}:=\frac{1}{2}\left\{\frac{\partial g(x,y)}{\partial x}+j\frac{\partial g(x,y)}{\partial y}\right\}.$
B-B Complex Hessian Matrices
The Hessian matrices to real-valued functions $g(\boldsymbol{z})$ with their argument $\boldsymbol{z}\in\mathbb{C}^{T}$ are defined in terms of (Wirtinger) complex gradient operations relative to the variables $\boldsymbol{z}$ and $\boldsymbol{z}^{*}$ as follows
$H_{c}(\boldsymbol{z}):=\left[\begin{array}[]{c|c}H_{\boldsymbol{z}^{*}\boldsymbol{z}}&\left(H_{\boldsymbol{z}^{\top}\boldsymbol{z}}\right)^{*}\\ \hline\cr H_{\boldsymbol{z}^{\top}\boldsymbol{z}}&\left(H_{\boldsymbol{z}^{*}\boldsymbol{z}}\right)^{\top}\end{array}\right],$ (57)
where matrices $H_{\boldsymbol{z}^{*}\boldsymbol{z}},H_{\boldsymbol{z}^{T}\boldsymbol{z}}$ correspond to the operations
$H_{\boldsymbol{z}^{*}\boldsymbol{z}}:=\nabla_{\boldsymbol{z}^{*}}\left[\nabla_{\boldsymbol{z}}g(\boldsymbol{z})\right],$ (58)
$H_{\boldsymbol{z}^{\top}\boldsymbol{z}}:=\nabla_{\boldsymbol{z}^{\top}}\left[\nabla_{\boldsymbol{z}}g(\boldsymbol{z})\right]$
and
$\nabla_{\boldsymbol{z}^{\top}}g(\boldsymbol{z})=\left[\nabla_{\boldsymbol{z}}g(\boldsymbol{z})\right]^{\top}.$ (59)
Appendix C Properties of the Function in (14)
In this section, we demonstrate that the function $H(x,{\boldsymbol{y}})$ in (14) satisfies Assumption 1. It is evident that $H(x,{\boldsymbol{y}})\geq 0$.
Notice that
$\nabla_{x}H(x,{\boldsymbol{y}})=$ $-2\Re\{{\boldsymbol{y}^{H}}\boldsymbol{E}^{H}{\frac{d\boldsymbol{a}}{dx}}\},$ (60)
$\nabla^{2}_{xx}H(x,{\boldsymbol{y}})=$ $-2\Re\{{\boldsymbol{y}^{H}}\boldsymbol{E}^{H}{\frac{d^{2}\boldsymbol{a}}{dx^{2}}}\}$
$=$ $-2\Re\{{\boldsymbol{y}^{H}}\boldsymbol{E}^{H}{\boldsymbol{D}^{2}}\boldsymbol{a}(x)\}.$
where ${\frac{d\boldsymbol{a}}{dx}}=\boldsymbol{D}\boldsymbol{a}(x)$ and $\boldsymbol{D}={\operatorname{Diag}}(0,j,2j,\cdots,(N-1)j)$. Thus,
$|\nabla^{2}_{xx}H(x_{1},{\boldsymbol{y}})-\nabla^{2}_{xx}H(x_{2},{\boldsymbol{y}})|$ (61)
$=2\Re\{{\boldsymbol{y}^{H}}\boldsymbol{E}^{H}{\boldsymbol{D}^{2}}(\boldsymbol{a}(x_{1})-\boldsymbol{a}(x_{2}))\}$
$\leq 2\|{\boldsymbol{D}^{2}}\boldsymbol{E}\boldsymbol{{y}}\|\|\boldsymbol{a}(x_{1})-\boldsymbol{a}(x_{2})\|.$
Let $g(x)=\frac{1}{jk}e^{jkx}$. Since $|\frac{d^{2}}{dx^{2}}g(x)|\leq k$, $g(x)$ is Lipschitz continuous with moduli $L=k$. Therefore,
$|\frac{d}{dx}g(x_{1})-\frac{d}{dx}g(x_{2})|$ $=|e^{jkx_{1}}-e^{jkx_{2}}|$
$\leq k|x_{1}-x_{2}|,\forall x_{1},x_{2}\in\mathbb{R}.$
Then,
$\|\boldsymbol{a}(x_{1})-\boldsymbol{a}(x_{2})\|^{2}$ $=\frac{1}{N}\sum_{k=1}^{N-1}(e^{jkx_{1}}-e^{jkx_{2}})^{2}$ (62)
$\leq\frac{1}{N}\sum_{k=1}^{N-1}k^{2}(x_{1}-x_{2})^{2}$
$=\frac{(N-1)(2N-3)}{6}(x_{1}-x_{2})^{2}.$
From (61) and (62), we obtain
$|\nabla^{2}_{xx}H(x_{1},{\boldsymbol{y}})-\nabla^{2}_{xx}H(x_{2},{\boldsymbol{y}})|\leq$ $L(\boldsymbol{{y}})|x_{1}-x_{2}|,$ (63)
$\forall x_{1},x_{2}\in\mathbb{R},$
where $L(\boldsymbol{{y}})=\sqrt{\frac{2(N-1)(2N-3)}{3}}\|{\boldsymbol{D}^{2}}\boldsymbol{E}\boldsymbol{{y}}\|$. This indicates that the function $x\rightarrow H(x,\boldsymbol{y})$ is $C_{L(\boldsymbol{y})}^{2,2}$.
For any fixed $x$, we have
${(\nabla^{2}_{\boldsymbol{y}\boldsymbol{y}})_{c}H(x,\boldsymbol{y})}={\boldsymbol{I}_{2T}}\succ{\boldsymbol{0}},$
which indicates that the function $\boldsymbol{y}\rightarrow H(x,\boldsymbol{y})$ is $\nu$-strongly convex with $\nu=2$.
According to [24], consider a function $H(x,\boldsymbol{y})$ whose partial Hessian $\nabla^{2}_{xx}H(x,\boldsymbol{y})$ is Lipschitz continuous. Let $\mu^{-}$ be an arbitrary positive constant, and replace the Lipschitz modulus $L(\boldsymbol{y})$ by $L^{\prime}(\boldsymbol{y})=\max\{L(\boldsymbol{y}),\mu\}$. Then, $L^{\prime}(\boldsymbol{y})$ remains a Lipschitz modulus for $\nabla^{2}_{xx}H(x,\boldsymbol{y})$. Consequently, the inequality in (12) is trivially satisfied with $L^{\prime}(\boldsymbol{y})$ and $\lambda^{-}=\mu^{-}$. Under the assumption that the sequence $\{\boldsymbol{z}^{k}\}_{k\in\mathbb{N}}$ generated from the proposed algorithm is bounded, the inequality in (13) holds because $L(\boldsymbol{{y}})$ is upper-bounded by $\|\boldsymbol{{y}}\|$ with constant scaling.